If x4+y4=16, use the following steps yo find y''.

a. Use implicit differentiation to find y'.

b. Use the Quotient Rule to differentiate the expression for y' from part (a). Express your answer in terms of x and y only.


1
Expert's answer
2021-11-16T11:11:09-0500

a.

"4x^3+4y^3y'=0"

"y'=-x^3\/y^3"


b.

"y''=-\frac{3x^2y^3-3y^2x^3y'}{y^6}=-\frac{3x^2y^3+3x^6/y}{y^6}=-\frac{3x^2y^4+3x^6}{y^7}"